OpenStudy (anonymous):

PLEASE!!!!!!!!!!!!!!!! A Project to restore a house, a painter used 8 5/12 gallons of primer, 10 1/2 gallons of paint and 6 1/2 of enamel. Find the total number of gallons of paint used?

OpenStudy (anonymous):

8 + 10 + 6 = convert the fractions to 12ths 5/12 + 6/12 + 6/12 = 17/12 = 1 5/12 add this to the sum of the whole numbers above